# RankMixer: Scaling Up Ranking Models in Industrial Recommenders

### v1

@@ Abstract.
Recent progress on large language models (LLMs) has spurred interest in scaling up recommendation systems, yet two practical obstacles remain. First, training and serving cost on industrial Recommenders must respect strict latency bounds and high QPS demands. Second, most human-designed feature-crossing modules in ranking models were inherited from the CPU era and fail to exploit modern GPUs, resulting in low Model Flops Utilization (MFU) and poor scalability. We introduce RankMixer, a hardware-aware model design tailored towards a unified and scalable feature-interaction architecture. RankMixer retains the transformer’s high parallelism while replacing quadratic self-attention with multi-head token mixing module for higher efficiency. Besides, RankMixer maintains both the modeling for distinct feature subspaces and cross-feature-space interactions with Per-token FFNs. We further extend it to one billion parameters with a Sparse-MoE variant for higher ROI. A dynamic routing strategy is adapted to address the inadequacy and imbalance of experts training. Experiments show RankMixer’s superior scaling abilities on a trillion-scale production dataset. By replacing previously diverse handcrafted low-MFU modules with RankMixer, we boost the model MFU from 4.5% to 45%, and scale our ranking model parameters by 100x while maintaining roughly the same inference latency. We verify RankMixer’s universality with online A/B tests across three core application scenarios (Recommendation, Advertisement and Search). Finally, we launch 1B Dense-Parameters RankMixer for full traffic serving without increasing the serving cost, which improves user active days by 0.2% and total in-app usage duration by 0.5%.
@@ 1. Introduction
3. To address these challenges, we propose a hardware-aware model design method, RankMixer. The core design of RankMixer is based on two scalable components: 1. Multi-head token mixing achieves cross-token feature interactions only through the parameter-free operator. This strategy outperforms the self-attention mechanism in terms of performance and computational efficiency. 2. Per-token feed-forward networks (FFNs) expand model capacity significantly and tackle inter-feature-space domination problem by allocating independent parameters for different feature subspace modeling. These FFNs also align well with the recommendation data patterns, enabling a better scaling behavior. To further boost the ROI of large-scale models, we extend the per-token FFN modules into a Sparse Mixture-of-Experts (MoE) structure. By dynamically activating only a subset of experts per token for different data, we can significantly increase the model capacity with a minimal increase in computational cost. RankMixer adopts a highly parallel architecture similar to transformers, but overcomes several critical limitations of self-attention based feature-interaction: low training efficiency, the combinatorial explosion when modeling cross-space ID similarity and severe memory-bound caused by attention weights matrix. At the same time, RankMixer offers greater model capacity and learning capability under the same Flops compared with Vanilla Transformer.
@@ 3.4. Sparse MoE in RankMixer
To further increase the scaling ROI, we replace dense FFNs with Sparse Mixture-of-Experts (MoE) blocks, so that the model’s capacity grows while computation cost remains roughly constant. Vanilla Sparse-MoE, however, degrades in RankMixer because: (i) uniform $k$-expert routing. Top-$k$ selection treats all feature tokens equally, wasting the budget on low-information tokens and starving high-information ones, which hinders the model from capturing differences between tokens. (ii) expert under-training. Per-token FFNs already multiply parameters by #tokens; adding non-shared experts further explodes the expert count, producing highly unbalanced routing and poorly trained experts;
@@ 4.1.2. Evaluation Metrics
To evaluate model performance, we use AUC (Area Under the Curve) and UAUC (User-level AUC) as the primary performance metrics and Parameter count, FLOPs and MFU as the efficiency metrics listed as follows: Finish/Skip AUC/UAUC: A finish=1/0 or skip=1/0 label indicates whether a user completed watching a video or slide to the next video in a short-time. We evaluate the AUC and UAUC for this finish label. an AUC increase of 0.0001 can be regarded as a confidently significant improvement. Dense-Param: The number of parameters in the dense part, excluding the sparse embedding parameters. Training Flops/Batch: The number of floating-point operations (FLOPs) required to run a single training batch through the model, representing the computational cost of training. The batch size is set as 1024. MFU: MFU(Model FLOPs Utilization) is a metric that measures how effectively a model utilizes floating-point operations provided by the hardware, calculated by dividing the model’s actual FLOPs consumption by the hardware’s theoretical FLOPs capacity.
@@ 4.1.3. Baselines
All experiments were conducted on hundreds of GPUs in a hybrid distributed training framework that the sparse part is updated asynchronously, while the dense part is updated synchronously. The optimizer hyperparameters were kept consistent across all models. For the dense part, we used the RMSProp optimizer with a learning rate of 0.01, while the sparse part used the Adagrad optimizer. The batch size was set to 1024.
@@ 4.5. Sparse-MoE Scalability and Expert Balance
Table 5. Online lift of RankMixer in Advertising and Search scenarios
Advertising Search
@@
Lift +0.73% +3.90% +1.75% +0.1414% $-1.0$%
@@ 4.6. Online Serving cost
Table 6. Case Study on Model Efficiency
Model #Param Flops GFlops/#Param(M) MFU Latency
Base-DLRM-8.7M 8.7M 52G 5.9 4.51% 16.12ms
Wukong(l=8,nL=32) 122M 442G 3.6 18.51% 33.7ms
RankMixer-1B 1B 2106 G 2.1 44.57% 14.3ms
@@ 4.7. Online Performance
To verify the universality of RankMixer as a scaling recommendation model framework, we conducted online experiments in the three core application scenarios of personalised ranking—feed recommendation, advertising, and search—thus covering all major use-cases of personalized ranking. For each scenario we monitor the following key performance indicators:
@@
[…]
The Previous baselines in these scenarios are 10 Million-level parameter DLRM model. We replaced the Dense part with the RankMixer-1B model. We conduct online A/B test experiments across Feed Recommendation, Advertising and Search. The A/B test result on Feed Recommendation are shown on the Table 4 and the result of Advertising and Search Ranking are shown on the Table 5.
RankMixer was deployed and evaluated in the three personalised-ranking application including the Feed Recommendation (RankMixer-1B), Advertising (RankMixer-1B), and in-app Search (RankMixer-300M). RankMixer delivered statistically significant uplifts on all business-critical metrics. We can also observe that the improvements on low-active users are largest on the Table 4 compared to other activeness level user groups with over 0.45% Active Days improvements, which demonstrates the model’s strong generalization capability. These results demonstrate that the RankMixer as a unified backbone generalises reliably to different application across all important scenarios.

### v2

@@ Abstract.
Recent progress on large language models (LLMs) has spurred interest in scaling up recommendation systems, yet two practical obstacles remain. First, training and serving cost on industrial Recommenders must respect strict latency bounds and high QPS demands. Second, most human-designed feature-crossing modules in ranking models were inherited from the CPU era and fail to exploit modern GPUs, resulting in low Model Flops Utilization (MFU) and poor scalability. We introduce RankMixer, a hardware-aware model design tailored towards a unified and scalable feature-interaction architecture. RankMixer retains the transformer’s high parallelism while replacing quadratic self-attention with multi-head token mixing module for higher efficiency. Besides, RankMixer maintains both the modeling for distinct feature subspaces and cross-feature-space interactions with Per-token FFNs. We further extend it to one billion parameters with a Sparse-MoE variant for higher ROI. A dynamic routing strategy is adapted to address the inadequacy and imbalance of experts training. Experiments show RankMixer’s superior scaling abilities on a trillion-scale production dataset. By replacing previously diverse handcrafted low-MFU modules with RankMixer, we boost the model MFU from 4.5% to 45%, and scale our online ranking model parameters by 70x while maintaining roughly the same inference latency. We verify RankMixer’s universality with online A/B tests across two core application scenarios (Recommendation and Advertisement). Finally, we launch 1B Dense-Parameters RankMixer for full traffic serving without increasing the serving cost, which improves user active days by 0.3% and total in-app usage duration by 1.08%.
@@ 1. Introduction
To address these challenges, we propose a hardware-aware model design method, RankMixer. The core design of RankMixer is based on two scalable components: 1. Multi-head token mixing achieves cross-token feature interactions only through the parameter-free operator. This strategy outperforms the self-attention mechanism in terms of performance and computational efficiency. 2. Per-token feed-forward networks (FFNs) expand model capacity significantly and tackle inter-feature-space domination problem by allocating independent parameters for different feature subspace modeling. These FFNs also align well with the recommendation data patterns, enabling a better scaling behavior. To further boost the ROI of large-scale models, we extend the per-token FFN modules into a Sparse Mixture-of-Experts (MoE) structure. By dynamically activating only a subset of experts per token for different data, we can significantly increase the model capacity with a minimal increase in computational cost. RankMixer adopts a highly parallel architecture similar to transformers, but overcomes several critical limitations of self-attention based feature-interaction: low training efficiency, the combinatorial explosion when modeling cross-space ID similarity and severe memory-bound caused by attention weights matrix. At the same time, RankMixer offers greater model capacity and learning capability under the same Flops compared with Vanilla Transformer.
@@ 3.4. Sparse MoE in RankMixer
To further increase the scaling ROI, we can replace the dense FFNs of each pertoken with Sparse Mixture-of-Experts (MoE) blocks, so that the model’s capacity grows while computation cost remains roughly constant. Vanilla Sparse-MoE, however, degrades in RankMixer because: (i) uniform $k$-expert routing. Top-$k$ selection treats all feature tokens equally, wasting the budget on low-information tokens and starving high-information ones, which hinders the model from capturing differences between tokens. (ii) expert under-training. Per-token FFNs already multiply parameters by #tokens; adding non-shared experts further explodes the expert count, producing highly unbalanced routing and poorly trained experts;
@@ 4.1.2. Evaluation Metrics
To evaluate model performance, we use AUC (Area Under the Curve) and UAUC (User-level AUC) as the primary performance metrics and Parameter count, FLOPs and MFU as the efficiency metrics listed as follows: Finish/Skip AUC/UAUC: A finish=1/0 or skip=1/0 label indicates whether a user completed watching a video or slide to the next video in a short-time. We evaluate the AUC and UAUC for this finish label. an AUC increase of 0.0001 can be regarded as a confidently significant improvement. Dense-Param: The number of parameters in the dense part, excluding the sparse embedding parameters. Training Flops/Batch: The number of floating-point operations (FLOPs) required to run a single batch of 512 through the model, representing the computational cost of training. MFU: MFU(Model FLOPs Utilization) is a metric that measures how effectively a model utilizes floating-point operations provided by the hardware, calculated by dividing the model’s actual FLOPs consumption by the hardware’s theoretical FLOPs capacity.
@@ 4.1.3. Baselines
All experiments were conducted on hundreds of GPUs in a hybrid distributed training framework that the sparse part is updated asynchronously, while the dense part is updated synchronously. The optimizer hyperparameters were kept consistent across all models. For the dense part, we used the RMSProp optimizer with a learning rate of 0.01, while the sparse part used the Adagrad optimizer.
@@ 4.5. Sparse-MoE Scalability and Expert Balance
Table 5. Online lift of RankMixer in Advertising
Advertising
@@
Lift +0.73% +3.90%
Table 6. Metrics of Online Model Deployment and Cost
Metric Base-16M RankMixer-1B Change
# Dense Param 15.8M 1.1B $\uparrow$ 70$\times$
Flops 107G 2106G $\uparrow$ 20.7$\times$
Flops/Param(G/M) 6.8 1.9 $\downarrow$ 3.6$\times$
MFU 4.47% 44.57% $\uparrow$ 10$\times$
Inference Precisioin fp32 fp16 $\uparrow$ 2$\times$
Latency 14.5ms 14.3ms -
@@ 4.7. Online Performance
To verify the universality of RankMixer as a scaling recommendation model framework, we conducted online experiments in the two core application scenarios of personalised ranking—feed recommendation and advertising, covering major use-cases of personalized ranking. For each scenario we monitor the following key performance indicators:
@@
The Previous baselines in these scenarios are 10 Million-level parameter DLRM model. We replaced the Dense part with the RankMixer-1B model. We conduct online A/B test experiments across Feed Recommendation and Advertising. The A/B test result on Feed Recommendation are shown on the Table 4 and the result of Advertising is shown on the Table 5.
RankMixer was deployed and evaluated in the three personalised-ranking application including the Feed Recommendation (RankMixer-1B) and Advertising (RankMixer-1B). RankMixer delivered statistically significant uplifts on all business-critical metrics. We can also observe that the improvements on low-active users are largest on the Table 4 compared to other activeness level user groups with over 0.45% Active Days improvements, which demonstrates the model’s strong generalization capability. These results demonstrate that the RankMixer as a unified backbone generalises reliably to different application across all important scenarios.